A 0.08 m3 container has 5,854 molecules of salt A membrane of 0:007m thickness is connecting this container to an outside same volume container filled with pure water Assume that the diffusion coefficient of salt is 133 x 10-9 m2/s.


Calculate the initial concentration of salt in units of million molecules/m3


1
Expert's answer
2021-04-12T07:02:29-0400
"n_i=\frac{N_i}{V}\\n_i=\frac{5854}{0.08}=73175\frac{1}{m^3}=0.073175\frac{10^6}{m^3}"